Three capacitor are connected first in series and then in parallel. Find the equivalent capacitance
for each type of combination

Grade:Upto college level

1 Answers

Apoorva Arora IIT Roorkee
9 years ago
When the capacitors are connected in series then the equivalent capacitance is given by
\frac{1}{C_{eq}}=\frac{1}{C_{1}}+\frac{1}{C_{2}}+\frac{1}{C_{3}}

However, when the capacitors are connected in parallel, the equivalent capacitance is given by
C_{eq}=C_{1}+C_{2}+C_{3}